Suppose T € L(R^2) is defined by T(x;y) = ((3y; x). Find the eigenvalues of T.


1
Expert's answer
2021-07-12T17:42:42-0400

Let us solve the equation "T(x,y)=(\lambda x, \lambda y)" for "\lambda" :

"\begin{cases} 3y=\lambda x \\ x=\lambda y \end{cases}" , now using the substitution we find that "\begin{cases} 3y=\lambda^2y \\ x= \lambda^2x/3 \end{cases}".

Therefore, the possible solutions for "\lambda" are "\sqrt{3}, -\sqrt{3}" (as at least one of numbers "x,y" is non-zero).

Another way to calculate it could be expressing "T" in a matrix form :

"T= \begin{pmatrix} 0 & 3 \\ 1 & 0 \end{pmatrix}"

and find the characteristic polynomial :

"\lambda^2-3=0"

which yields the same result.